OpenStudy (anonymous):

Which expression is a difference of cubes? x^6-6 x^6-8 x^8-6 x^8-8

OpenStudy (anonymous):

This isnt math

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

XD

OpenStudy (welshfella):

what is it then?

OpenStudy (anonymous):

im pretty sure it is...

OpenStudy (anonymous):

NVM This isnt english XD XD XDDXDXDXDXD

OpenStudy (anonymous):

Im sorry buddy. Hope someone helps you out :)

OpenStudy (welshfella):

you have to be careful with this question Remember when you multiply terms with exponents like x^2 you ADD the exponents.

OpenStudy (welshfella):

example x^2 * x^2 = x^(2+2) = x^4

OpenStudy (anonymous):

yes i understand that bit but all and all im not understanding how to do this question

OpenStudy (welshfella):

which of the numbers is a perfect cube 6 or 8 ?

OpenStudy (anonymous):

6

OpenStudy (welshfella):

no

OpenStudy (welshfella):

2*2*2 is not 6

OpenStudy (anonymous):

okay i see what you're saying now. then its 8, so how do i determine the rest?

OpenStudy (welshfella):

You are now left with 2 choices x^6 or x^8 you now need to think of a term in x which when multiply by itself 3 times = x^6 or x^8

OpenStudy (welshfella):

- this where where the rule of exponents comes in...

OpenStudy (anonymous):

the answer should be x^6-8 right?

OpenStudy (welshfella):

right because x^2 * x^2 * x^2 = x^6

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

or written another way (x^2)^3 = x^(2*3) = x^6

OpenStudy (welshfella):

yw